Pool10 Move Status Thread.


08.21.2013 9:30PM PST

Nothing to report, all systems go. yet we are reporting it, so.. meta.

08.21.2013 4:30PM PST

New storage is running as smooth as silk. Engineers instructed not to blink, eat, or sleep for the next 12 hours to watch its every move. Sorry again for the trouble.

08.21.2013 4:15PM PST

New storage devices in place, monitoring performance and health.

08.21.2013 3:25PM PST

Switching now.

08.21.2013 3:20PM PST

Preparing to switch to new storage blocks.

08.21.2013 11:30AM PST

Restarting storage. Implementing alternative storage solutions.

08.20.2013 7:28AM PST

Investigating storage slowdown.

08.17.2013 7:28AM PST

Services restarting.

08.16.2013 8:08AM PST

Normal Service restored.

08.16.2013 8:00AM PST

Services ready, will begin routing traffic back to pool momentarily.

08.16.2013 7:30AM PST

Network protocol to the primary storage block is experiencing issues.

08.16.2013 5:30AM PST

Team is working on restoring service.

08.15.2013 3:00PM PST

Unresolved Sites using timthumb are still being investigated for performance issues.

08.15.2013 1:00PM PST

Sites using timthumb are experiencing slowdowns, we are investigating.

08.15.2013 8:40AM PST

Traffic coming thru our UK nodes are properly resolving to pool10.

08.15.2013 8:30AM PST

A handful of requests routed through our UK facility are not resolving properly to the new pool. Correcting now.

08.15.2013 7:10AM PST

Investigating a handful of site not resolving correctly.

08.15.2013 1:10AM PST

Resuming normal operations.

08.15.2013 12:38AM PST

Routing traffic here temporarily while we make a change to the db master configs.

08.14.2013 10:40PM PST

Database cluster has stabilized. All systems operating normally and well within spec.

08.14.2013 10:10PM PST

Adding additional database capacity.

08.14.2013 10:00PM PST

Some database slowdown on the slaves, Engineers working on it.

08.14.2013 9:45PM PST

Traffic is now resolving to pool10, please contact support with any issues. We are correcting a few anomalies as we see them.

08.14.2013 9PM PST

Running final checks before switching traffic.

08.14.2013 8PM PST

We are delaying slightly before switching the traffic over.

08.14.2013 7PM PST

Engineers will be making the sync of the databases and final system prior to redirecting all traffic.

08.14.2013 5PM PST

Earlier today we sent the following email to approx 200 customers.  This blog post will track the status of the migration.

screenshot888

Pagely® is the original Managed WordPress Hosting company. Since 2009 we have helped thousands of personal bloggers, small business, and enterprise clients secure and accelerate their WordPress powered sites. Let us help you.

The #1 rated WordPress Host - WPMU.org

Give Us a Try Today